### v4

# Approximating Nash Social Welfare

Abstract
For any $\varepsilon>0$, we give a simple, deterministic $(4+\varepsilon)$-approximation algorithm for the Nash social welfare (NSW) problem under submodular valuations. We also consider the asymmetric variant of the problem, where the objective is to maximize the weighted geometric mean of agents’ valuations, and give an $\mathrm{e}(\omega+2+\varepsilon)$-approximation if the ratio between the largest weight and the average weight is at most $\omega$.
We also show that the $\nicefrac{{1}}{{2}}$-EFX envy-freeness property can be attained simultaneously with a constant-factor approximation. More precisely, we can find an allocation in polynomial time that is both $\nicefrac{{1}}{{2}}$-EFX and a $(8+\varepsilon)$-approximation to the symmetric NSW problem under submodular valuations.
1 Introduction
We consider the problem of allocating a set $G$ of $m$ indivisible items among a set $A$ of $n$ agents, where each agent $i\in A$ has a valuation function $v_{i}:2^{G}\to{\mathbb{R}}_{\geq 0}$ and a weight (entitlement) $w_{i}>0$ such that $\sum_{i\in A}w_{i}=1$. The Nash social welfare (NSW) problem asks for an allocation ${\mathcal{S}}=(S_{i})_{i\in A}$ that maximizes the weighted geometric mean of the agents’ valuations,
$\operatorname*{NSW}({\mathcal{S}})=\prod_{i\in A}\left(v_{i}(S_{i})\right)^{w_{i}}.$
We refer to the special case when all agents have equal weight (i.e., $w_{i}=1/n$) as the symmetric NSW problem, and call the general case the asymmetric NSW problem. Throughout, we let $w_{\max}\coloneqq\max_{i\in A}w_{i}$. For $\alpha>1$, an $\alpha$-approximate solution to the NSW problem is an allocation $\mathcal{S}$ with $\operatorname*{NSW}(\mathcal{S})\geq\mathrm{OPT}/\alpha$, where $\mathrm{OPT}$ denotes the optimum value of the NSW-maximization problem.
Allocating resources among agents in a fair and efficient manner is a fundamental problem in computer science, economics, and social choice theory; we refer the reader to the monographs [5, 12, 13, 50, 55, 56, 58] on the background. A common measure of efficiency is utilitarian social welfare, i.e., the sum of the utilities $\sum_{i\in A}v_{i}(S_{i})$ for an allocation $(S_{i})_{i\in A}$. In contrast, fairness is often measured by max-min fairness, i.e., $\min_{i\in A}v_{i}(S_{i})$; maximizing this objective is also known as the Santa Claus problem [4].
Symmetric NSW provides a balanced tradeoff between the often conflicting requirements of fairness and efficiency. It has been introduced independently in a variety of contexts. It is a discrete analogue of the Nash bargaining game [42, 51]; it corresponds to the notion of competitive equilibrium with equal incomes in economics [57]; and arises as a proportional fairness notion in networking [43]. The more general asymmetric objective has also been well-studied since the seventies [40, 41]. It has found applications in different areas, such as bargaining theory [20, 44], water resource allocation [27, 39], and climate agreements [59].
A distinctive feature of the NSW problem is invariance under scaling of the valuation functions $v_{i}$ by independent factors $\lambda_{i}$, i.e., each agent can express their preference in a “different currency” without changing the optimization problem (see [50] for additional characteristics).
Submodular and subadditive valuation functions
A set function $v:\,2^{G}\to{\mathbb{R}}$ is monotone if $v(S)\leq v(T)$ whenever $S\subseteq T$. A monotone set function with $v(\emptyset)=0$ is also called a valuation function or simply valuation. The function $v:\,2^{G}\to{\mathbb{R}}$ is submodular if
$v(S)+v(T)\geq v(S\cap T)+v(S\cup T)\,\quad\forall S,T\subseteq G\,$,
and subadditive if
$v(S)+v(T)\geq v(S\cup T)\,\quad\forall S,T\subseteq G\,.$
We assume the valuation functions are given by value oracles that return $v(S)$ for any $S\subseteq G$ in $O(1)$ time.
$\nicefrac{{1}}{{2}}$-EFX allocations
Envy-freeness is one of the most desirable fair division concepts. However, for indivisible goods, exact envy-freeness is generally unattainable: if there is only one desirable item, all but one agent will necessarily be envious. This limitation has motivated a rich line of work introducing and analyzing various relaxations. Among these, envy-freeness up to any item (EFX) is the strongest and widely regarded as the most compelling fairness notion in the discrete setting with equal entitlements [19]. An allocation $\mathcal{S}=(S_{i})_{i\in A}$ is said to be EFX if
$v_{i}(S_{i})\geq v_{i}(S_{k}\setminus\{j\}),\enspace\forall i,k\in A\,,\forall j\in S_{k}\ .$
That is, no agent envies another agent’s bundle after the removal of any single item from that agent’s bundle. It is not known whether EFX allocations always exist or not, even in the additive case with more than three agents, and it is regarded as “fair division’s biggest open question” [54]. This motivated the study of its relaxation $\alpha$-EFX for an $\alpha\in(0,1)$, where an allocation $\mathcal{S}$ is said to be $\alpha$-EFX if
$v_{i}(S_{i})\geq\alpha\cdot v_{i}(S_{k}\setminus\{j\}),\enspace\ \forall i,k\in A\,,\forall j\in S_{k}\,.$
The best-known $\alpha$, for which the existence is known, is $\nicefrac{{1}}{{2}}$ for submodular valuations, albeit with the efficiency guarantee of $O(n)$-approximation to the symmetric NSW problem [53, 23].
Our contributions
Our main theorem on NSW is the following.
Theorem 1.1.
For any $\varepsilon>0$, there is a deterministic polynomial-time $(nw_{\max}+2+\varepsilon)\mathrm{e}$-approximation algorithm for the asymmetric Nash social welfare problem with submodular valuations. For symmetric instances, the algorithm returns a $(4+\varepsilon)$-approximation. The number of arithmetic operations and value oracle calls is polynomial in $n$, $m$, and $1/\varepsilon$.
Algorithm 1 in Section 2 presents the algorithm asserted in the theorem. Note that $nw_{\max}$ is the ratio between the maximum weight $w_{\max}$ and the average weight ($1/n$). In the symmetric case, when all weights are $w_{i}={1}/{n}$, this bound gives $(3+\varepsilon)\mathrm{e}<8.2$. In this case, we can improve the analysis to obtain a $(4+\varepsilon)$-approximation algorithm.
As our second result, we show that a $\nicefrac{{1}}{{2}}$-EFX allocation with high NSW value exists and can also be efficiently found. We give a general reduction for subadditive valuations. An allocation $\mathcal{S}=(S_{i})_{i\in A}$ is called complete if every item is included, i.e., $\bigcup_{i\in A}S_{i}=G$; otherwise, it is called partial. In the context of $\nicefrac{{1}}{{2}}$-EFX allocations, $\operatorname*{NSW}(\mathcal{S})$ will always refer to the NSW value of allocation $\mathcal{S}$ in the symmetric case ($w_{i}=1/n$ for all $i\in A$).
Theorem 1.2.
There is a deterministic strongly polynomial-time algorithm that given a symmetric NSW instance with subadditive valuations and given a (complete or partial) allocation $\mathcal{S}$ of the items, it returns a complete allocation $\mathcal{T}$ that is $\nicefrac{{1}}{{2}}$-EFX and $\operatorname*{NSW}(\mathcal{T})\geq\operatorname*{NSW}(\mathcal{S})/2$.
The above algorithm is strongly polynomial in the value oracle model: the number of basic arithmetic operations and oracle calls is polynomially bounded in $n$ and $m$. Together with Theorem 1.1, we obtain the following corollary.
Corollary 1.3.
For any $\varepsilon>0$, there is a deterministic polynomial algorithm that returns a $\nicefrac{{1}}{{2}}$-EFX complete allocation that is $(8+\varepsilon)$-approximation to the symmetric NSW problem under submodular valuations. The number of arithmetic operations and value oracle calls is polynomial in $n$, $m$, and $1/\varepsilon$.
1.1 Related work
Prior work on approximating NSW
Let us first consider additive valuations, i.e., when $v_{i}(S)=\sum_{j\in S}v_{ij}$ for nonnegative values $v_{ij}$. Maximizing symmetric NSW is NP-hard already in the case of two agents with identical additive valuations, by a reduction from the Subset-Sum problem. It is NP-hard to approximate within a factor better than $1.069$ for additive valuations [35], and better than $1.5819$ for submodular valuations [38].
On the positive side, a number of remarkably different constant-factor approximations are known for additive valuations. The first such algorithm with the factor of $2\cdot e^{1/e}\approx 2.889$ was given by Cole and Gkatzelis [26] using a continuous relaxation based on a particular market equilibrium concept. Later, [25] improved the analysis of this algorithm to achieve the factor of 2. Anari, Oveis Gharan, Saberi, and Singh [2] used a convex relaxation that relies on properties of real stable polynomials. The current best factor is $\mathrm{e}^{1/\mathrm{e}}+\varepsilon\simeq 1.45$ by Barman, Krishnamurthy, and Vaish [8]; the algorithm uses a different market equilibrium based approach and also guarantees EF1 fairness, a weaker version of the EFX property.
For the general class of subadditive valuations [6, 23], $O(n)$-approximations are known. This is the best one can hope for in the value oracle model [6], for the same reasons that this is impossible to improve for the utilitarian social welfare problem [29]. Sublinear $O(n^{53/54})$ approximation was designed for XOS valuations if we are given access to both demand and XOS oracles [7]. Recall that all submodular valuations are XOS, and all XOS valuations are subadditive.
Constant-factor approximations were also obtained beyond additive valuation functions: capped-additive [34], separable piecewise-linear concave (SPLC) [3], and their common generalization, capped-SPLC [21] valuations; the approximation factor for capped-SPLC valuations matches the $\mathrm{e}^{1/\mathrm{e}}+\varepsilon$ factor for additive valuations. All these valuations are special classes of submodular. Subsequently, Li and Vondrák [46] designed an algorithm that estimates the optimal value within a factor of $\frac{\mathrm{e}^{3}}{(\mathrm{e}-1)^{2}}\simeq 6.8$ for a broad class of submodular valuations, such as coverage and summations of matroid rank functions, by extending the techniques of [2] using real stable polynomials. However, this algorithm only estimates the optimum value but does not find a corresponding allocation in polynomial time.
In [37], Garg, Husić, and Végh developed a constant-factor approximation for a subclass of submodular valuations called Rado-valuations. These include weighted matroid rank functions and many others that can be obtained using operations such as induction by network and contractions. An important example outside this class is the coverage valuation. They attained an approximation ratio 772 for the symmetric case and $772(w_{\max}/w_{\min})^{3}$ for the asymmetric case. Most recently, Li and Vondrák [47] obtained a randomized 380-approximation for symmetric NSW under submodular valuations by extending the the approach of [37].
We significantly improve and simplify the approach used in [37] and [47]; we give a comparison to these works in Section 2.5.
Subsequent work on approximating NSW
Subsequent to the conference version of this work [36], Dobzinski, Li, Rubinstein, and Vondrák [28] obtained a constant factor approximation for symmetric NSW for subadditive valuations, assuming a demand oracle. They extend the approach of [37] and [47], using also a configuration LP rounding.
Further, Brown, Laddha, Pittu, and Singh [14] obtained an approximation factor of $5\exp(2\log n+2\sum_{i\in A}w_{i}\log w_{i})$ for additive valuations. The term in $\exp(.)$ is twice the Bregman divergence between the distribution $(w_{i})_{i\in A}$ and the uniform distribution. As pointed out by Bento Natura [52], a simple change in our analysis leads to the improved bound $\exp(\log(3n)+\sum_{i\in A}w_{i}\log w_{i})\cdot(\mathrm{e}+\varepsilon)$ in Theorem 1.1 for general submodular valuations; this is explained in Remark 2.15. Feng and Li [32] obtained an $(\mathrm{e}^{1/\mathrm{e}}+\varepsilon)$-approximation for general asymmetric NSW with additive valuations using a novel rounding of the configuration LP. Soon after, their approach was generalized by Feng, Hu, Li, and Zhang [31] to get a $(233+\varepsilon)$-approximation algorithm for the asymmetric NSW with general submodular valuations. Most recently, Bei, Feng, Hu, Li, and Zhang [10] further improved this ratio to $5.18$. Moreover, [10] also tightened the analysis of the algorithm in Theorem 1.1, improving the factor $4+\varepsilon$ to $3.914+\varepsilon$.
Prior work on EFX and related notions
The existence of EFX allocations has not been settled despite significant efforts [19, 53, 22, 54]. This problem is open for more than two agents with general monotone valuations (including submodular), and for more than three agents with additive valuations. This necessitated the study of its relaxations $\alpha$-EFX for $\alpha\in(0,1)$ and partial EFX allocations. For the notion of $\alpha$-EFX, the best-known $\alpha$ is $0.618$ for additive [1] and $0.5$ for subadditive valuations (including submodular) [53].
For the notion of partial EFX allocations, the existence is known for general monotone valuations if we do not allocate at most n2n-2 items [24, 49, 11], albeit without any efficiency guarantees. For additive valuations, although n2n-2 is still the best bound known, there exist partial EFX allocations with 2-approximation to the NSW problem [18].
A well-studied weaker notion is envy-freeness up to one item (EF1), where no agent envies another agent after the removal of some item from the envied agent’s bundle. EF1 allocations are known to exist for general monotone valuations and can also be computed in polynomial-time [48]. However, an EF1 allocation alone is not desirable because it might be highly inefficient in terms of any welfare objective. For additive valuations, the allocations maximizing NSW are EF1 [19]. Although the NSW problem is APX-hard [45], there exists a pseduopolynomial time algorithm to find an allocation that is EF1 and 1.45-approximation to the NSW problem under additive valuations [8]. For capped-SPLC valuations, [21] shows the existence of an allocation that is $\nicefrac{{1}}{{2}}$-EF1 and $1.45$-approximation to the NSW problem.
Subsequent to the conference version of this work, Feldman, Mauras, and Pontika [30] improved Theorem 1.2 to show the existence of an allocation $\mathcal{T}$ that is $\nicefrac{{1}}{{2}}$-EFX and $\operatorname*{NSW}(\mathcal{T})\geq\nicefrac{{2}}{{3}}\operatorname*{NSW}(\mathcal{S})$ for a given allocation $\mathcal{S}$. For subadditive valuations, Barman and Suzuki [9] recently showed the existence of a complete allocation that is EF1 and provides a 2-approximation to the NSW problem, as well as of a partial allocation that is EFX and also provides a 2-approximation to the NSW problem.
Notation
We use $\log(x)$ for the natural logarithm throughout. For set $S\subseteq G$ and $j\in G$, we use $S+j$ to denote $S\cup\{j\}$ and SjS-j for $S\setminus\{j\}$ and we write $v(j)$ for $v(\{j\})$. For a vector $p\in{\mathbb{R}}^{G}$ and $S\subseteq G$, we denote $p(S)=\sum_{i\in S}p_{i}$.
By a matching from $A$ to $G$ we mean a mapping $\tau:\,A\to G\cup\{\bot\}$ where $\tau(i)\neq\tau(j)$ if $\tau(i)\neq\bot$; $\bot$ is a special symbol representing unmatched agents. For $X\subseteq A$, a matching is $X$-perfect, if $\tau(i)\neq\bot$ for every $i\in X$. An $A$-perfect matching is simply a perfect matching when every node on both sides is matched, which is possible when $|A|=|G|$.
Recall that the definition of valuation functions requires $v(\emptyset)=0$. In our algorithm, we will use the concept of endowed valuation functions. The set function $v\,:\,2^{G}\to\mathbb{R}$ is an endowed valuation function, if it is monotone, and $2v(\emptyset)\geq\max_{g\in G}v(g)$. Intutively, this means that the initial value is already at least the value of any singleton.
Overview of the paper
Section 2 describes the NSW approximation algorithm. A comparison of our technique with previous approaches is given in Section 2.5. The exposition of the fairness guarantee is given in Section 3. We conclude the paper with some remarks in Section 4.
2 Approximation algorithm for Nash social welfare
The NSW algorithm asserted in Theorem 1.1 is shown in Algorithm 1. The subroutine LocalSearch will be described in Algorithm˜2. Throughout, we assume $0<\varepsilon\leq 1$; we can always replace $\varepsilon$ by $\min\{\varepsilon,1\}$. The algorithm proceeds in the following three phases.
Input: Valuations $(v_{i})_{i\in A}$ over $G$, weights $w\in\mathbb{R}_{>0}^{A}$ such that $\sum_{i\in A}w_{i}=1$, and $\varepsilon>0$.
Output: Allocation $\mathcal{S}=(S_{i})_{i\in A}$.
1 Find an $A$-perfect matching $\tau:A\to G$ maximizing $\prod_{i\in A}(v_{i}(\tau(i)))^{w_{i}}$ and set $H\coloneqq\tau(A),J\coloneqq G\setminus H,\bar{A}:=\{i\in A:\,v_{i}(J)>0\}$
2 for $i\in\bar{A}$ do
3 $\ell(i)\leftarrow\operatorname*{argmax}\{v_{i}(\ell):\ell\in J\}$
4   Define $\bar{v}_{i}(S)\coloneqq v_{i}(\ell(i))+v_{i}(S)$
[…]
6 Find a matching $\sigma:A\to H$ maximizing $\prod_{i\in A}(v_{i}(R_{i}+\sigma(i)))^{w_{i}}$
7 return $\mathcal{S}=(R_{i}+\sigma(i))_{i\in A}$
Algorithm 1 Approximating the submodular NSW problem
Phase 1: Initial matching
We find an optimal assignment of one item to each agent, i.e., a matching $\tau:A\to G$ maximizing $\prod_{i\in A}(v_{i}(\tau(i)))^{w_{i}}$. This can be done using a maximum-weight matching algorithm with weights $w_{i}\log v_{i}(j)$ in the bipartite graph between $A$ and $G$ with edge set $\{(i,j):\,v_{i}(j)>0\}$. If no matching of cardinality $n=|A|$ exists, then we can conclude that there is no allocation with positive NSW value, and return an arbitrary allocation. For the rest of the paper, we assume there is a matching covering $A$, and let $H\coloneqq\tau(A)$ be the set of matched items.
Phase 2: Local search
In the second phase, we let $J\coloneqq G\setminus H$ denote the set of items not assigned in the first phase. Recall that we let $\bar{A}:=\{i\in A:\,v_{i}(J)>0\}$ denote the set of agents that have a positive value for some item in $J$. For every $i\in A$, we select
$\ell(i)\in\operatorname*{argmax}_{j\in J}v_{i}(j)$
as a favorite item of agent $i$ in $J$ (for agents in $A\setminus\bar{A}$, this will be an arbitrary item with $v_{i}(\ell(i))=0$.) By submodularity, $v_{i}(\ell(i))>0$. For each $i\in\bar{A}$, we define the endowed valuation function $\bar{v}_{i}:2^{J}\to{\mathbb{R}}_{>0}$ as
$\bar{v}_{i}(S)\coloneqq v_{i}(\ell(i))+v_{i}(S)\quad\forall S\subseteq J\,.$
Thus, $\bar{v}_{i}(\emptyset)=v_{i}(\ell(i))$, and $\bar{v}_{i}(j)\leq 2\bar{v}_{i}(\emptyset)$ for any $j\in J$. Further, we set the accuracy parameter
$\bar{\varepsilon}\coloneqq\frac{\varepsilon}{2m}\,.$ (1)
Note that by the assumption $\varepsilon\leq 1$, the following bound holds:
$(1+\bar{\varepsilon})^{m}<1+\varepsilon\,.$ (2)
Our local search starts with allocating all items to a single agent in $\bar{A}$. As long as moving one item to a different agent increases the potential function
$\prod_{i\in\bar{A}}\left(\bar{v}_{i}(R_{i})\right)^{w_{i}}$
by at least a factor $(1+\bar{\varepsilon})$, we perform such an exchange. Phase 2 terminates when no more such exchanges are possible, and returns the current allocation. For all agents $i\in A\setminus\bar{A}$, we let $R_{i}=\emptyset$.
1 $R_{k}\leftarrow J$ for some $k\in\bar{A}$ and $R_{i}\leftarrow\emptyset$ for $i\in A\setminus\{k$}
2 while $\exists i,k\in\bar{A}$ and $j\in R_{i}$ such that $\left(\frac{\bar{v}_{i}(R_{i}-j)}{\bar{v}_{i}(R_{i})}\right)^{w_{i}}\cdot\left(\frac{\bar{v}_{k}(R_{k}+j)}{\bar{v}_{k}(R_{k})}\right)^{w_{k}}>1+\bar{\varepsilon}$ do
3 $R_{i}\leftarrow R_{i}-j$ and $R_{k}\leftarrow R_{k}+j$
return $\mathcal{R}:=(R_{i})_{i\in A}$
[…]
Phase 3: Rematching
In the final phase, we match the items in $H$ to the agents optimally, considering allocation $\mathcal{R}=(R_{i})_{i\in A}$ of $J$. This can be done by again solving a maximum-weight matching problem, now with weights $w_{ij}=w_{i}\log v_{i}(R_{i}+j)$.
In the remainder of this section, we prove Theorem 1.1. In Section 2.1, presents the main ‘Rematching Lemma’ (Lemma 2.1) that asserts that Algorithm 1 has value at least corresponding to the following solution: for every agent, take the best of $R_{i}$, $\ell(i)$, and $\pi(i)$, where $\pi(i)$ is a matching assigning to every agent their most valuable item from $H$ in an optimal solution. Note that $\max\{v_{i}(R_{i}),v_{i}(\ell(i))\}\geq\frac{1}{2}\bar{v}_{i}(R_{i})$, hence, this enables us to relate the value of the solution to the local search with respect to the endowed valuations.
In Section 2.2, we formulate simple properties of approximate local optimal solution found in Phase 2. We then compare the allocation of items in $J$ in an optimal solution to $(R_{i})_{i\in A}$ via a marginal pricing argument. This is done separately for the symmetric case (Section 2.3) and the asymmetric case (Section 2.4). The arguments follow the same lines, but the marginal prices are defined differently in the two settings. These together complete the proof of Theorem 1.1.
2.1 The Rematching Lemma
A crucial step in our algorithm relates the value of the assignment output by Algorithm 1 to the contribution of $H$ to an optimal allocation. Suppose that the optimal solution is $(H_{i}\cup S_{i})_{i\in A}$, where $H_{i}\subseteq H$ and $S_{i}\subseteq J$. This optimal solution does not necessarily assign $H$ via a matching—i.e., the cardinalities of $H_{i}$ could be arbitrary. Given $(H_{i})_{i\in A}$, we define a matching $\pi:A\to H$ as follows: If $H_{i}\neq\emptyset$, let $\pi(i)\coloneqq\arg\max_{j\in H_{i}}\bar{v}_{i}(j)$. If $H_{i}=\emptyset$, then define $\pi(i)$ arbitrarily so that $\pi$ is a perfect matching between $A$ and $H$.
A key lemma is the following. We note that the argument corresponds to a more general statement about bipartite matchings which might be of independent interest. We present and discuss this statement in Appendix A.
Lemma 2.1 (Rematching Lemma).
The value obtained by Algorithm 1 is at least
$\prod_{i\in A}\left(\max\{v_{i}(\pi(i)),v_{i}(\ell(i)),v_{i}(R_{i})\}\right)^{w_{i}}.$
Proof.
The allocation returned by our algorithm is $(R_{1}+\sigma(1),R_{2}+\sigma(2),\ldots,R_{n}+\sigma(n))$ where $(R_{1},\ldots,R_{n})$ is the partition of $X\setminus H$ found by the local search, and $\sigma:A\to H$ is the matching found in the last stage of our algorithm. Instead, let us consider the following allocation, which is more restricted: Given $H$ and $(R_{1},\ldots,R_{n})$, each agent can choose either their set $R_{i}$, or one of the items in $H$, but not both. We prove that the best such allocation satisfies the guarantee of the lemma.
Without loss of generality, let us denote the set of agents as $A=\{1,2,\ldots,n\}$, with $\bar{A}=\{1,2,\ldots,\bar{n}\}$ for $\bar{n}\leq n$. Let us construct the following bipartite graph: one side includes $A$, and the other side has $n+2\bar{n}$ nodes, labeled $\{h_{1},\ldots,h_{n},\ell_{1},\ldots,\ell_{\bar{n}},r_{1},\ldots,r_{\bar{n}}\}$. The nodes $h_{1},\ldots,h_{n}$ represent the elements of $H$. The nodes $\ell_{1},\ldots,\ell_{\bar{n}}$ represent the items $\ell(1),\ldots,\ell(\bar{n})$ — however, $\ell_{1},\ldots,\ell_{\bar{n}}$ are distinct nodes even if some items $\ell(i)$ coincide. Finally, the nodes $r_{1},\ldots,r_{\bar{n}}$ represent the sets $R_{1},\ldots,R_{\bar{n}}$. We define an edge $(i,h_{j})$ for each $i\in A$ and $h_{j}\in H$ with $v_{i}(h_{j})>0$, of weight $w_{i}\log v_{i}(h_{j})$. We also define an edge $(i,\ell_{i})$ of weight $w_{i}\log v_{i}(\ell(i))$, and an edge $(i,s_{i})$ of weight $w_{i}\log v_{i}(R_{i})$ for every $i\in\bar{A}$.
Figure 1: Proof of Lemma 2.1, illustration: the alternating path $h_{3}-3-h_{4}-4-\ell_{4}$ can be used to improve either $\rho$ or $\tau$.
[…]
Let $\rho$ be a maximum-weight matching of cardinality $|A|$ in this bipartite graph. Clearly, its weight is at least $\sum_{i=1}^{n}w_{i}\log\max\{v_{i}(\pi(i)),v_{i}(\ell(i)),v_{i}(R_{i})\}$, because one possible matching is to choose for each $i\in A$ the edge corresponding to the maximum of the 3 quantities. An issue with this matching is that it does not correspond to a feasible allocation, due to possible overlaps between the items $\ell(i)$ and the sets $R_{i}$.
The key claim is that there is a maximum-weight matching in this graph which does not use any of the edges $(i,\ell_{i})$. We prove this as follows: Suppose that $\rho$ is a maximum-weight matching that uses as few of the $(i,\ell_{i})$ edges as possible. Recall also the matching $\tau$ constructed in the initial stage, which is the maximum-weight matching between $A$ and the full item set $X$. In our graph, $\tau$ matches $A$ to $H=\{h_{1},\ldots,h_{n}\}$. Suppose that $\rho$ uses an edge $(i^{*},\ell_{i^{*}})$ and consider the alternating path $P$ in $\rho\Delta\tau$, starting from this edge ($\tau$ does not match the node $\ell_{i^{*}}$, so this is indeed the start of an alternating path). Except for the node $\ell_{i^{*}}$, the alternating path hops between $A$ and $H$, because starting from $\ell^{*}$, an edge of $\rho$ takes us to $A$, then an edge of $\tau$ takes us to $H$, etc. Eventually, the alternating path terminates at a node of $H$ which is not matched by $\rho$ (whereas each node of $A$ is matched by both $\tau$ and $\rho$). Now, consider two cases.
(1) The weight of the $\tau$-edges on this alternating path $P$ is at least as large as the weight of the $\rho$-edges on $P$. Then we can perform a swap, $\rho^{\prime}=\rho\Delta P$. This defines a matching $\rho^{\prime}$ in our graph whose weight is at least as large as that of $\rho$, and the number of $(i,\ell_{i})$ edges has decreased—a contradiction.
(2) The weight of the $\tau$-edges on $P$ is strictly less than the weight of the $\rho$-edges on $P$. Then we can perform the swap $\tau^{\prime}=\tau\Delta P$. Note that $\tau^{\prime}$ is a valid matching of items to agents, because it uses only $H$ and one item $\ell(i^{*})$, which is outside of $H$, so no conflict can arise here. Thus we obtain a matching $\tau^{\prime}:A\to X$ of weight strictly larger than $\tau$, which means $\tau$ was not a maximum-weight matching in the initial stage—a contradiction.
We conclude that there is a maximum-weight matching in our graph which does not use any $(i,\ell_{i})$ edge, and hence it corresponds to a feasible allocation of either $R_{i}$ or an element of $H$ to each agent. ∎
2.2 Characterizing Local Optima
In order to upper bound the optimal NSW value to the value obtained in Algorithm 1 via Lemma 2.1, we also need to relate $S_{i}$, the part of the optimal solution in $J$, to $R_{i}$ and $\ell(i)$. Note that Lemma 2.1 does not use the local optimality of $R_{i}$ but works for an arbitrary partition of $J$. In this subsection, we focus on properties of locally optimal solutions.
Recall that we work with the item set $J$, set of agents $\bar{A}$, favourite items $\ell(i)$, and endowed valuations $\bar{v}_{i}(S)=v_{i}(\ell(i))+v_{i}(S)$.
Definition 2.2 ($\bar{\varepsilon}$-local optimum).
A complete allocation $\mathcal{R}=(R_{i})_{i\in A}$ is an $\bar{\varepsilon}$-local optimum with respect to valuations $\bar{v}_{i}$, if for all pairs of different agents $i,k\in\bar{A}$ and all $j\in R_{i}$ it holds that
$\left(\frac{\bar{v}_{i}(R_{i}-j)}{\bar{v}_{i}(R_{i})}\right)^{w_{i}}\cdot\left(\frac{\bar{v}_{k}(R_{k}+j)}{\bar{v}_{k}(R_{k})}\right)^{w_{k}}\leq 1+\bar{\varepsilon}\,.$
A 0-local optimum will be simply called a local optimum.
Lemma 2.3.
[…]
Proof.
It is immediate that if the algorithm terminates, it returns an $\bar{\varepsilon}$-local maximum. Recalling that $\bar{v}_{i}(j)\leq 2\bar{v}_{i}(\emptyset)$ for any $j\in J$, submodularity implies $v_{i}(J)<(|J|+1)\bar{v}_{i}(\emptyset)\leq m\bar{v}_{i}(\emptyset)$ for every $i\in\bar{A}$, using also that $|J|\leq m-n\leq m-1$. Hence,
$\prod_{i\in\bar{A}}\left({\bar{v}_{i}(J})\right)^{w_{i}}\leq m\prod_{i\in\bar{A}}\left({\bar{v}_{i}(\emptyset)}\right)^{w_{i}}\,$,
and therefore the product $\prod_{i\in\bar{A}}{\bar{v}_{i}(R_{i})^{w_{i}}}$ may grow by at most a factor of $m$ throughout all exchange steps. Every swap increases this product by at least a factor $(1+\bar{\varepsilon})$. Thus, the total number of swaps is bounded by $\log_{1+\bar{\varepsilon}}m=m\log_{1+\varepsilon}m=O\left(\frac{m}{\varepsilon}\log m\right)$. ∎
We need the following two properties of submodular functions.
Proposition 2.4.
Let $\bar{v}:2^{J}\to{\mathbb{R}}_{>0}$ be a monotone submodular function. Let $S\subseteq T\subseteq J$ and $j\in J$. Then,
$\frac{\bar{v}(T+j)}{\bar{v}(T)}\leq\frac{\bar{v}(S+j)}{\bar{v}(S)}\,.$
Proof.
By the monotonicity and submodularity of $v$ we have
$\begin{aligned} \frac{\bar{v}(T+j)}{\bar{v}(T)}=\frac{\bar{v}(T)+\bar{v}(T+j)-\bar{v}(T)}{\bar{v}(T)}&\leq\frac{\bar{v}(S)+\bar{v}(T+j)-\bar{v}(T)}{\bar{v}(S)}\\ &\leq\frac{\bar{v}(S)+\bar{v}(S+j)-\bar{v}(S)}{\bar{v}(S)}=\frac{\bar{v}(S+j)}{\bar{v}(S)}.\end{aligned}\qed$
Proposition 2.5.
Let $\bar{v}_{i}:2^{J}\to{\mathbb{R}}_{>0}$ be the submodular endowed valuation for an agent $i\in\bar{A}$. For any $j\in R$,
$\bar{v}_{i}(R-j)\geq\sum_{k\in R}(\bar{v}_{i}(R)-\bar{v}_{i}(R-k)).$
Proof.
Let us denote $R-j\coloneqq\{r_{1},\dots,r_{s}\}$. By submodularity, we have
$\bar{v}_{i}(R-j)$ $=\bar{v}_{i}(\emptyset)+\sum_{k=1}^{s}(\bar{v}_{i}(\{r_{1},\dots,r_{k}\})-\bar{v}_{i}(\{r_{1},\dots,r_{k-1}\}))$
$\geq\bar{v}_{i}(\emptyset)+\sum_{k=1}^{s}(\bar{v}_{i}(R)-\bar{v}_{i}(R-r_{k}))\geq\sum_{k\in R}(\bar{v}_{i}(R)-\bar{v}_{i}(R-r_{k}))$
where in the last step, we used the fact that $\bar{v}_{i}(\emptyset)=v_{i}(\ell(i))\geq v_{i}(j)\geq\bar{v}_{i}(R)-\bar{v}_{i}(R-j)$ . ∎
We analyze our local search in slightly different ways in the symmetric case and the general asymmetric case. We consider the symmetric case first.
2.3 Local optimality analysis for symmetric NSW
Recall that $\bar{\varepsilon}=\varepsilon/(2m)$. We define another accuracy parameter for the analysis as
$\hat{\varepsilon}\coloneqq(1+\bar{\varepsilon})^{n}-1\,.$ (3)
Thus, $1+\hat{\varepsilon}=(1+\bar{\varepsilon})^{n}\leq(1+\bar{\varepsilon})^{m}=1+\varepsilon$ from (2) and $n\leq m$. In what follows, we complete the proof of Theorem 1.1 for the symmetric case, i.e., $w_{1}=\ldots=w_{n}=1/n$, showing the approximation factor $4+\varepsilon$.
Let $j\in J$ and let $i\in\bar{A}$ be the agent such that $j\in R_{i}$. We define the price of $j$ as
$p_{j}\coloneqq\frac{\bar{v}_{i}(R_{i})}{\bar{v}_{i}(R_{i}-j)}-1=\frac{\bar{v}_{i}(R_{i})-\bar{v}_{i}(R_{i}-j)}{\bar{v}_{i}(R_{i}-j)}\,.$
The following lemma shows that the spending of each agent $i$, $p(R_{i})=\sum_{j\in R_{i}}p_{j}$, is at most $1$.
Lemma 2.6 (Bounded spending).
Let $\mathcal{R}=(R_{i})_{i\in A}$ be an $\bar{\varepsilon}$-local optimum with respect to the endowed valuations $\bar{v}_{i}$. Then, $p(R_{i})\leq 1$ for every agent $i\in\bar{A}$, and consequently, $p(J)\leq|\bar{A}|\leq n$.
Proof.
From the definition of the prices $p_{j}$, and by Proposition 2.5, we have
$p(R_{i})=\sum_{j\in R_{i}}\frac{\bar{v}_{i}(R_{i})-\bar{v}_{i}(R_{i}-j)}{\bar{v}_{i}(R_{i}-j)}\leq\frac{\sum_{j\in R_{i}}(\bar{v}_{i}(R_{i})-\bar{v}_{i}(R_{i}-j))}{\min_{k\in R_{i}}\bar{v}_{i}(R_{i}-k)}\leq 1\,.$
Since $(R_{1},\ldots,R_{n})$ is a partition of $J$ (every item is allocated throughout the local search), we have
$p(J)=\sum_{j\in J}p_{j}=\sum_{i\in\bar{A}}\sum_{j\in R_{i}}p_{j}\leq|\bar{A}|\,.\qed$
The next lemma shows that the prices can be used to bound the marginal increase in utility by adding new items to the $R_{k}$ bundle of an agent.
Lemma 2.7.
Given an $\bar{\varepsilon}$-local optimum $\mathcal{R}=(R_{i})_{i\in\bar{A}}$, and the prices $p_{j}$ defined as above, for every $k\in\bar{A}$ and $T\subseteq J$, it holds that
$\frac{\bar{v}_{k}(R_{k}\cup T)}{\bar{v}_{k}(R_{k})}\leq 1+p(T)+2{\hat{\varepsilon}}|T|\,.$
Proof.
First, note that
$p_{j}\leq 1\quad\forall j\in J\,.$ (4)
This follows already by Lemma 2.6; one can also see it directly since by construction of $\bar{v}_{i}$ and submodularity, $\bar{v}_{i}(R_{i})-\bar{v}_{i}(R_{i}-j)\leq v_{i}(j)\leq\bar{v}_{i}(\emptyset)\leq\bar{v}_{i}(R_{i}-j)$. Hence, $p_{j}=\frac{(\bar{v}_{i}(R_{i})-\bar{v}_{i}(R_{i}-j))}{\bar{v}_{i}(R_{i}-j)}\leq 1$.
Let $i\in\bar{A}$ such that $j\in R_{i}$. From the $\bar{\varepsilon}$-optimality of $\mathcal{R}$, using also (2), and (4), we get
$\frac{\bar{v}_{k}(R_{k}+j)}{\bar{v}_{k}(R_{k})}\leq(1+\bar{\varepsilon})^{n}\frac{\bar{v}_{i}(R_{i})}{\bar{v}_{i}(R_{i}-j)}=(1+\hat{\varepsilon})(1+p_{j})\leq 1+p_{j}+2\hat{\varepsilon}\,$, (5)
because otherwise we could swap item $j$ to agent $k$; recall that $w_{k}=w_{i}=1/n$. Using submodularity together with this bound, we have
$\frac{\bar{v}_{k}(R_{k}\cup T)}{\bar{v}_{k}(R_{k})}$ $\leq\frac{\bar{v}_{k}(R_{k})+\sum_{j\in T}(\bar{v}_{k}(R_{k}+j)-\bar{v}_{k}(R_{k}))}{\bar{v}_{k}(R_{k})}$
$\leq 1+\sum_{j\in T}(p_{j}+2\hat{\varepsilon})\leq p(T)+1+2\hat{\varepsilon}|T|\,.$
As a small detour to illustrate the usefulness of the above bound, we show that local search gives a $2$-approximation of the optimal Nash social welfare with respect to the endowed valuations for the agents in $\bar{A}$ that have positive valuation for $J$.
Proposition 2.8.
Let $\mathcal{R}=(R_{i})_{i\in A}$ be a local optimum returned by Algorithm 2 and $\mathcal{S}=(S_{i})_{i\in A}$ be an optimal NSW allocation of $J$ with respect to the endowed submodular valuations $\bar{v}_{i}$. Then
$\left(\prod_{i\in\bar{A}}\bar{v}_{i}(R_{i})\right)^{1/n}\geq\frac{1}{2}\cdot\left(\prod_{i\in\bar{A}}\bar{v}_{i}(S_{i})\right)^{1/n}\,.$
Proof.
By Lemma 2.6, $\sum_{i\in\bar{A}}p(S_{i})=p(J)\leq n$. Then, by Lemma 2.7 and the AM-GM inequality,
$\prod_{i\in\bar{A}}\bar{v}_{i}(S_{i})$ $\leq\prod_{i\in\bar{A}}\bar{v}_{i}(R_{i}\cup S_{i})\leq\prod_{i\in\bar{A}}(1+p(S_{i}))\bar{v}_{i}(R_{i})$
$\leq\left(\frac{1}{n}\sum_{i\in\bar{A}}(1+p(S_{i}))\right)^{n}\prod_{i\in\bar{A}}\bar{v}_{i}(R_{i})$
$\leq 2^{n}\cdot\prod_{i\in\bar{A}}\bar{v}_{i}(R_{i})\,.\qed$
Proposition 2.8 is included solely for the intuition and we do not use it directly: we still need to take the items in $H$ into consideration and move from the endowed valuations $\bar{v}_{i}$ to the original valuations $v_{i}$. In what follows, we relate the locally optimal allocation of $J$, $(R_{i})_{i\in{A}}$ to the optimal solution $(H_{i}\cup S_{i})_{i\in A}$.
Lemma 2.9.
Let $\mathcal{R}=(R_{i})_{i\in A}$ be an $\bar{\varepsilon}$-local optimum with respect to the endowed valuations $\bar{v}_{i}$, and let $i\in\bar{A}$. For any set $H_{i}\subseteq H$, $S_{i}\subseteq J$,
$\bar{v}_{i}(R_{i}\cup S_{i}\cup H_{i})\leq(1+p(S_{i})+2\hat{\varepsilon}|S_{i}|)\,\bar{v}_{i}(R_{i})+|H_{i}|\,v_{i}(\pi(i))\,.$
Proof.
By submodularity and the choice of $\pi(i)$, $v_{i}(H_{i})\leq\sum_{j\in H_{i}}v_{i}(j)\leq|H_{i}|v_{i}(\pi(i))$. By Lemma 2.7 and submodularity again, the lemma follows. ∎
We are ready to prove the approximation guarantee in the symmetric case.
Lemma 2.10.
Let $\varepsilon\geq 0$, $\bar{\varepsilon}=\varepsilon/(2m)$, and $\mathcal{R}=(R_{i})_{i\in A}$ be an $\bar{\varepsilon}$-local optimum with respect to the endowed valuations $\bar{v}_{i}$. Then, there exists a matching $\rho:A\to H$ such that
$\operatorname*{NSW}(\mathcal{R},\rho)\geq\frac{\operatorname{OPT}}{4(1+\varepsilon)}\,.$
Proof.
Consider the optimal allocation $(H_{i}\cup S_{i})_{i\in A}$ as above. Denote
$V_{i}\coloneqq\max\{v_{i}(\pi(i)),v_{i}(\ell(i)),v_{i}(R_{i})\}\,$,
and recall that according to Lemma 2.1, our algorithm achieves value at least $\left(\prod_{i=1}^{n}V_{i}\right)^{1/n}$. First, consider an agent $i\in\bar{A}$. By Lemma 2.9,
$\bar{v}_{i}(S_{i}\cup H_{i})\leq\bar{v}_{i}(R_{i}\cup S_{i}\cup H_{i})\leq(1+p(S_{i})+2\hat{\varepsilon}|S_{i}|)\,\bar{v}_{i}(R_{i})+|H_{i}|v_{i}(\pi(i)).$
Equivalently, since $\bar{v}_{i}(S)=v_{i}(S)+v_{i}(\ell(i))$,
$v_{i}(S_{i}\cup H_{i})$ $\leq v_{i}(R_{i})+(p(S_{i})+2\hat{\varepsilon}|S_{i}|)(v_{i}(R_{i})+v_{i}(\ell(i)))+|H_{i}|v_{i}(\pi(i))$
$\leq(1+2p(S_{i})+4\hat{\varepsilon}|S_{i}|+|H_{i}|)V_{i}\,.$
The same bound also holds for $i\in A\setminus\bar{A}$, noting that $v_{i}(S_{i}\cup H_{i})=v_{i}(H_{i})\leq|H_{i}|v_{i}(\pi(i))$. Hence, by Lemma 2.1, we get
$\operatorname{OPT}=\left(\prod_{i=1}^{n}v_{i}(S_{i}\cup H_{i})\right)^{1/n}\leq\prod_{i=1}^{n}\left((1+2p(S_{i})+4\hat{\varepsilon}|S_{i}|+|H_{i}|)V_{i}\right)^{1/n}.$
By the AM-GM inequality,
$\operatorname{OPT}$ $\leq$ $\prod_{i=1}^{n}\left(1+2p(S_{i})+4\hat{\varepsilon}|S_{i}|+|H_{i}|\right)^{1/n}\left(\prod_{i=1}^{n}V_{i}\right)^{1/n}$
$\leq$ $\left(\frac{1}{n}\sum_{i=1}^{n}\left(1+2p(S_{i})+4\hat{\varepsilon}|S_{i}|+|H_{i}|\right)\right)\left(\prod_{i=1}^{n}V_{i}\right)^{1/n}$
$\leq$ $4\left(1+\frac{m\hat{\varepsilon}}{n}\right)\left(\prod_{i=1}^{n}V_{i}\right)^{1/n}$
using the facts that $\sum_{i=1}^{n}p(S_{i})\leq n$, $\sum_{i=1}^{n}|H_{i}|\leq n$ and $|S_{i}|\leq m$. The proof is complete since $\left(1+\frac{m\hat{\varepsilon}}{n}\right)\leq(1+\hat{\varepsilon})^{m/n}=(1+\bar{\varepsilon})^{m}<(1+\varepsilon)$ by the definition of $\hat{\varepsilon}$ and by (2). ∎
2.4 Analysis for asymmetric NSW
We now modify the above analysis to the case of general weights $w_{i}$. The main difference is in the defintion of the prices $p_{j}$. For $\bar{\varepsilon}=\varepsilon/(2m)$, let $\mathcal{R}=(R_{i})_{i\in A}$ be an $\bar{\varepsilon}$-local optimum with respect to the endowed valuations $\bar{v}_{i}$. Let $j\in J$ and let $i\in\bar{A}$ be the agent such that $j\in R_{i}$. We define the price of $j$ as
$p_{j}\coloneqq w_{i}\log\frac{\bar{v}_{i}(R_{i})}{\bar{v}_{i}(R_{i}-j)}\,.$
We first show the analogue of Lemma 2.6. Note that the total spending is now bounded as $p(J)\leq 1$ rather than $p(J)\leq n$; this is mainly due to the incorporation of the weights $w_{i}$ in the definition of the prices.
Lemma 2.11 (Bounded spending).
For an $\bar{\varepsilon}$-local optimum $\mathcal{R}=(R_{i})_{i\in\bar{A}}$ and prices $p_{j}$ defined as above, $p(R_{i})\leq w_{i}$ for every agent $i\in\bar{A}$, and consequently, $p(J)\leq 1$.
Proof.
From the definition of $p_{j}$, we have
$p(R_{i})=w_{i}\sum_{j\in R_{i}}\log\frac{\bar{v}_{i}(R_{i})}{\bar{v}_{i}(R_{i}-j)}\leq w_{i}\sum_{j\in R_{i}}\frac{\bar{v}_{i}(R_{i})-\bar{v}_{i}(R_{i}-j)}{\bar{v}_{i}(R_{i}-j)}\leq w_{i}$
due to the elementary inequality $\log x\leq x-1$, and by Proposition 2.5 we know that $\sum_{j\in R_{i}}(\bar{v}_{i}(R_{i})-\bar{v}_{i}(R_{i}-j))\leq\bar{v}_{i}(R_{i}-{j^{\prime}})$ for $j^{\prime}\in\operatorname*{argmin}_{j\in R_{i}}\bar{v}_{i}(R_{i}-{j})$.
Adding up the prices over all the sets $R_{i}$, whose union is $J$, we obtain $p(J)=\sum_{i\in\bar{A}}p(R_{i})\leq\sum_{i\in\bar{A}}w_{i}\leq 1$. ∎
Lemma 2.12.
For endowed valuations $\bar{v}_{k}$, an $\bar{\varepsilon}$-local optimum $\mathcal{R}=(R_{i})_{i\in\bar{A}}$ and prices $p_{j}$ defined as above, for every item $j\in R_{i}$ and every agent $k\in\bar{A}$, we have
$\frac{\bar{v}_{k}(R_{k}+j)}{\bar{v}_{k}(R_{k})}\leq(1+\bar{\varepsilon})^{1/w_{k}}\mathrm{e}^{p_{j}/w_{k}}\,.$
Moreover, if the endowed valuation $\bar{v}_{k}$ is submodular, for all $T\subseteq J$, we have
$\frac{\bar{v}_{k}(R_{k}\cup T)}{\bar{v}_{k}(R_{k})}\leq(1+\bar{\varepsilon})^{|T|/w_{k}}\cdot\mathrm{e}^{p(T)/w_{k}}\,.$
Proof.
By definition, $\mathrm{e}^{p_{j}/w_{i}}=\frac{\bar{v}_{i}(R_{i})}{\bar{v}_{i}(R_{i}-j)}$. If $k=i$ the first statement is trivial. Otherwise, for $k\neq i$, the first statement follows from the $\bar{\varepsilon}$-optimality of $\mathcal{R}$; if false, we would swap item $j$ to agent $k$.
For the second statement, let us denote $T=\{t_{1},t_{2},\ldots,t_{|T|}\}\subseteq J$. Since $\bar{v}_{k}$ is submodular, by Proposition 2.4 we have
$\begin{aligned} \frac{\bar{v}_{k}(R_{k}\cup T)}{\bar{v}_{k}(R_{k})}=\prod_{a=1}^{|T|}\frac{\bar{v}_{k}(R_{k}\cup\{t_{1},\dots,t_{a}\})}{\bar{v}_{k}(R_{k}\cup\{t_{1},\dots,t_{a-1}\})}&\leq\prod_{a=1}^{|T|}\frac{\bar{v}_{k}(R_{k}+t_{a})}{\bar{v}_{k}(R_{k})}\\ &\leq\left(1+\bar{\varepsilon}\right)^{|T|/w_{k}}\mathrm{e}^{p(T)/w_{k}}\,.\end{aligned}\qed$
Similarly to Lemma 2.9, we have the following lemma for the bound of optimum solutions using prices and values of singletons in $H$. Recall from Section 2.1 that $(S_{i}\cup H_{i})_{i\in A}$ is an optimal solution with $\pi(i)=\arg\max_{j\in H_{i}}v(j)$ if $H_{i}\neq\emptyset$.
Lemma 2.13.
Let $\mathcal{R}=(R_{i})_{i\in A}$ be an $\bar{\varepsilon}$-local optimum with respect to the endowed valuations $\bar{v}_{i}$. For $i\in\bar{A}$, any set $H_{i}\subseteq H$, $S_{i}\subseteq J$,
$\bar{v}_{i}(R_{i}\cup S_{i}\cup H_{i})\leq(1+\bar{\varepsilon})^{|S_{i}|/w_{i}}\mathrm{e}^{p(S_{i})/w_{i}}\bar{v}_{i}(R_{i})+|H_{i}|v_{i}(\pi(i)).$
Proof.
By submodularity and the choice of $\pi(i)$, $v_{i}(H_{i})\leq\sum_{j\in H_{i}}v_{i}(j)\leq|H_{i}|v_{i}(\pi(i))$. By Lemma 2.12 and submodularity again, the lemma follows. ∎
Our final approximation bound follows by combining this with Lemma 2.1.
Lemma 2.14.
Let $\varepsilon\geq 0$, $\bar{\varepsilon}=\varepsilon/(2m)$, and let $\mathcal{R}=(R_{i})_{i\in A}$ be an $\bar{\varepsilon}$-local optimum with respect to the endowed valuations $\bar{v}_{i}$. Then, there exists a matching $\rho:A\to H$ such that
$\operatorname*{NSW}(\mathcal{R},\rho)\geq\frac{\operatorname{OPT}}{(2+nw_{\max})\mathrm{e}(1+\varepsilon)}\,.$
Proof.
Let the optimal allocation to agent $i\in A$ be $S_{i}\cup H_{i}$, as above, with $\pi(i)=\operatorname*{argmax}_{j\in H_{i}}v(j)$. Let
$V_{i}\coloneqq\max\{v_{i}(\pi(i)),v_{i}(\ell(i)),v_{i}(R_{i})\}\,.$
Thus, Lemma 2.1 shows that Algorithm 1 returns a solution with value at least $\prod_{i\in A}V_{i}^{w_{i}}$. By Lemma 2.13, for $i\in\bar{A}$,
$\bar{v}_{i}(S_{i}\cup H_{i})\leq\bar{v}_{i}(R_{i}\cup S_{i}\cup H_{i})\leq(1+\bar{\varepsilon})^{|S_{i}|/w_{i}}\mathrm{e}^{p(S_{i})/w_{i}}\bar{v}_{i}(R_{i})+|H_{i}|v_{i}(\pi(i)),.$
Equivalently, since $\bar{v}_{i}(S)=v_{i}(S)+v_{i}(\ell(i))$,
$v_{i}(S_{i}\cup H_{i})$ $\leq v_{i}(R_{i})+((1+\bar{\varepsilon})^{|S_{i}|/w_{i}}\mathrm{e}^{p(S_{i})/w_{i}}-1)(v_{i}(R_{i})+v_{i}(\ell(i)))+|H_{i}|v_{i}(\pi(i))$
$\leq\left(2(1+\bar{\varepsilon})^{|S_{i}|/w_{i}}\mathrm{e}^{p(S_{i})/w_{i}}-1+|H_{i}|\right)V_{i}\,.$
The same bound trivially holds for agents $i\in A\setminus\bar{A}$. Hence,
$\operatorname{OPT}=\prod_{i=1}^{n}\left(v_{i}(S_{i}\cup H_{i})\right)^{w_{i}}\leq\prod_{i=1}^{n}\left(2(1+\bar{\varepsilon})^{|S_{i}|/w_{i}}e^{p(S_{i})/w_{i}}+|H_{i}|)V_{i}\right)^{w_{i}}.$
We now get
$\operatorname{OPT}$ $\leq$ $\prod_{i=1}^{n}\left(2(1+\bar{\varepsilon})^{|S_{i}|/w_{i}}\mathrm{e}^{p(S_{i})/w_{i}}+|H_{i}|\right)^{w_{i}}V_{i}^{w_{i}}$
$\leq$ $\prod_{i=1}^{n}\left(2+|H_{i}|\right)^{w_{i}}(1+\bar{\varepsilon})^{|S_{i}|}\mathrm{e}^{p(S_{i})}V_{i}^{w_{i}}$
$\leq$ $\left(\sum_{i=1}^{n}w_{i}(2+|H_{i}|)\right)(1+\bar{\varepsilon})^{m}\mathrm{e}\,\prod_{i=1}^{n}V_{i}^{w_{i}}$
$\leq$ $(2+nw_{\max})(1+\varepsilon)\mathrm{e}\,\prod_{i=1}^{n}V_{i}^{w_{i}}\,.$
Here, the second inequality follows since $(1+\bar{\varepsilon})^{|S_{i}|/w_{i}}\mathrm{e}^{p(S_{i})/w_{i}}\geq 1$, the third inequality follows by the AM-GM inequality and by $\sum_{i=1}^{n}p(S_{i})\leq 1$ (Lemma 2.11); the final inequality uses $\sum_{i=1}^{n}|H_{i}|\leq n$, and (2). ∎
Remark 2.15 (Natura [52]).
The bound in Theorem 1.1 can be in fact strengthened to $\exp\big{(}\log(3n)+\sum_{i=1}^{n}w_{i}\log w_{i}\big{)}\cdot(\mathrm{e}+\varepsilon)$, a slight improvement over the bound in [14] for the additive case. This follows by strengthening the bound $(1+\varepsilon)(2+nw_{\max})$ in Lemma 2.10 to $3n(1+\varepsilon)\prod_{i\in A}w_{i}^{w_{i}}$. To see this, we only need to change the final estimate on $\prod_{i=1}^{n}\left(2+|H_{i}|\right)^{w_{i}}$ in the proof. Here, we are given $h_{i}\geq 0$ such that $\sum_{i\in A}h_{i}\leq n$, and need to upper bound $\prod_{i\in A}(2+h_{i})^{w_{i}}$. Instead of simply using AM-GM, we use AM-GM for a reweighted product:
$\prod_{i\in A}(2+h_{i})^{w_{i}}=\prod_{i\in A}w_{i}^{w_{i}}\prod_{i\in A}\left(\frac{2+h_{i}}{w_{i}}\right)^{w_{i}}\leq\prod_{i\in A}w_{i}^{w_{i}}\sum_{i\in A}(2+h_{i})\leq 3n\prod_{i\in A}w_{i}^{w_{i}}\,.$
2.5 Our techniques and comparison with previous approaches
The ideas in this paper were gradually developed by a series of work by subsets of the authors, namely, the conference papers [37] and in [47], and the joint paper [36]. The present paper represents further simplifications of [36], and can be thus seen as the joint journal version of these three papers.
At a high level, all three algorithms proceed in three phases, with Phases 1 and 3 being the same as outlined above. However, they largely differ in how the allocation $\mathcal{R}$ of $J=G\setminus H$ is obtained in Phase 2.
Garg, Husić, and Végh [37] use a rational convex relaxation, based on the concave extension of Rado valuations. After solving the relaxation exactly, they use combinatorial arguments to sparsify the support of the solution and construct an integral allocation. However, this relaxation and rounding exploits the combinatorial structure of Rado valuations, in particular, that the relaxation with the concave extension can be solved exactly.
Li and Vondrák [47] obtained the first polynomial-time algorithm for arbitrary submodular valuations. A significant challenge is that for submodular functions, the concave extension is NP-hard to evaluate. Instead, they work with the multilinear extension, which can be evaluated with random sampling, but it is not concave. To solve the relaxation (approximately), they use an iterated continuous greedy algorithm. The allocation $\mathcal{R}$ is obtained by independent randomized rounding of this fractional solution. Whereas the algorithm is simple, the analysis is somewhat involved. The main tool to analyze the rounding is the Efron–Stein concentration inequality; but this only works well if every item in the support of the fractional solution has bounded value. This is not true in general, and the argument instead analyzes a two-stage randomized rounding that gives a lower bound on the performance of the actual algorithm. First, a set of “large” fractional items is preserved, and a careful combinatorial argument is needed to complete the allocation.
Our approach for the second part is radically different and much simpler. We do not use any continuous relaxation; instead, $\mathcal{R}$ is obtained by discrete local search. Our discrete local search works with a certain modification of the valuation functions, where each agent is endowed with their largest singleton $\ell(i)$ “for free”. Using these modified valuations, we can guarantee a high NSW value of an infeasible allocation $(R_{i}+\ell(i))_{i\in A}$ of $J$ in the analysis. Our pricing analysis of the local search is inspired by the conditional equilibrium notion introduced by Fu, Kleinberg, and Lavi [33]. Conditional equilibrium is a relaxation of Walrasian equilibrium of indivisible goods: this is an allocation and pricing where players cannot improve on their utility by adding new goods to their current bundle. They show that a conditional equilibrium attains at least half of the maximal utilitarian welfare; this corresponds to an approximate version of the first welfare theorem for quasilinear utilities. As shown in Proposition 2.8, our local search gives an analogous bound for Nash social welfare.
We note that local search applied directly to the NSW problem cannot yield a constant factor approximation algorithm even if we allow changing an arbitrary fixed number $k$ of items. This can be seen already when $m=n$, in which case every allocation with positive NSW value is a perfect matching. Also, some other natural variants of local search do not work, or the analysis is not clear; for example, our analysis does not seem to extend when local search is applied to the (seemingly more natural) choice of $\bar{v}_{i}(S)=v_{i}(S+\tau(i))$. The idea of defining $\ell(i)$ and using the modified valuation functions is inspired by rounding of the fractional solution from previous approaches; the role of the $\ell(i)$ items is similar to the large items in [47], but we obtain much better guarantees using a more direct deterministic approach. We remark that the iterated continuous greedy algorithm in [47] can be analogously replaced by continuous local search, with a slightly improved approximation factor—this variant is unpublished.
A key part of the analysis is the Rematching Lemma (Lemma 2.1) that shows how the value of the final solution $(R_{i}+\sigma(i))_{i\in A}$ can be related to the infeasible allocation $(R_{i}+\ell(i))_{i\in A}$. While the rematching phase was already present (and essentially identical) in [37] and [47], the analysis here is much simpler. At the heart of the argument is a general monotonicity property of maximum-weight bipartite matchings on subsets of nodes; we discuss this in Appendix A.
3 Finding fair and efficient allocations
In this section, we restate and prove Theorem 1.2. See 1.2 Our first key tool is a subroutine that finds a partial allocation that is $\nicefrac{{1}}{{2}}$-EFX and preserves a large fraction of the NSW value.
Lemma 3.1.
There exists a deterministic strongly polynomial algorithm MakeFairOrEfficient($\mathcal{T}$), that, for any partial allocation $\mathcal{T}$, returns another partial allocation $\mathcal{R}$ that satisfies one of the following properties
  1. (i)
    $\operatorname*{NSW}(\mathcal{R})\geq\operatorname*{NSW}(\mathcal{T})$ and $\cup_{i\in A}R_{i}\subsetneq\cup_{i\in A}T_{i}$, or
  2. (ii)
    $\operatorname*{NSW}(\mathcal{R})\geq\tfrac{1}{2}\operatorname*{NSW}(\mathcal{T})$ and $\mathcal{R}$ is $\nicefrac{{1}}{{2}}$-EFX.
This is shown by modifying the approach of Caragiannis, Gravin, and Huang [18]. The key subroutine for them provides a similar alternative as in Lemma 3.1. In outcome (ii), they have the stronger EFX guarantee, while in outcome (i), they show that the NSW value increases by a certain factor. In outcome (i), it is not clear how an increase in the NSW value could be shown for subadditive valuations. However, arguing about the support decrease leads to a simpler argument.
In [18], only a partial EFX allocation is found. Theorem 1.2 shows the existence of a complete allocation, albeit with the weaker $\nicefrac{{1}}{{2}}$-EFX property. To derive Theorem 1.2, we start by repeatedly calling MakeFairOrEfficient until outcome (ii) is reached. Note that the outcome (i) can only happen at most $m$ times because the number of items in $\mathcal{R}$ reduces by at least one after each call.
The allocation at this point may be partial. We show that the remaining items can be allocated using the classical envy-free cycle procedure by Lipton, Markakis, Mossel, and Saberi [48]. Even though this procedure is known for the weaker EF1 property [17], we show that—after a suitable preprocessing step—it can produce an $\nicefrac{{1}}{{2}}$-EFX allocation while not decreasing the NSW value of the allocation.
In the remainder of this section, we prove Theorem 1.2. In Section 3.1, we derive Lemma 3.1, and in Section 3.2, we derive Theorem 1.2 from this lemma.
3.1 Finding a fair or an efficient allocation
The subroutine MakeFairOrEfficient($\mathcal{T}$), shown in Algorithm 3, generalizes the algorithm by Caragiannis, Gravin, and Huang [18] from additive to subadditive valuations. We begin by defining the notions of $\nicefrac{{1}}{{2}}$-EFX feasible bundles and the associated feasibility graph.
Definition 3.2 ($\nicefrac{{1}}{{2}}$-EFX feasible bundles and graph).
Given a partial allocation $\mathcal{T}=(T_{i})_{i\in A}$, we say that $T_{k}$ is a $\nicefrac{{1}}{{2}}$-EFX feasible bundle for agent $i$, if
$v_{i}(T_{k})\geq\tfrac{1}{2}\max_{\ell\in A,j\in T_{\ell}}v_{i}(T_{\ell}-j)\,.$
The $\nicefrac{{1}}{{2}}$-EFX feasibility graph of $\mathcal{T}$ is a bipartite graph $\mathcal{K}=(A\cup\mathcal{T},E)$ where the edge set $E$ is defined as:
$E=$ $\left\{(i,T_{i})\ |\ T_{i}\text{ is }\text{$\nicefrac{{1}}{{2}}$-EFX feasible for }i\right\}\ \ \cup$ (6)
$\left\{(i,T_{k})\ |\ v_{i}(T_{k})>2v_{i}(T_{i})\text{ and }v_{i}(T_{k})\geq\max_{\ell\in A,j\in T_{\ell}}v_{i}(T_{\ell}-j)\right\}\,.$
The following claim follows directly from the definition.
Claim 3.3.
In the graph $\mathcal{K}=(A\cup\mathcal{T},E)$, every node $i\in A$ has degree at least 1.
Proof.
If the bundle $T_{i}$ is $\nicefrac{{1}}{{2}}$-EFX feasible for agent $i$, then by definition, the edge $(i,T_{i})$ is included in $E$, and thus the degree of $i$ is at least 1. Otherwise, there must exist a $k\in A$ and $j\in T_{k}$ such that $v_{i}(T_{k}-j)>2v_{i}(T_{i})$. Let $\ell\in A$ be the agent whose bundle maximizes $v_{i}(T_{\ell})$ among all agents. Then it follows that $v_{i}(T_{\ell})\geq v_{i}(T_{k}-j)>2v_{i}(T_{i})$, which implies, by the graph construction, that the edge $(i,T_{\ell})\in E$. Hence, every node $i\in A$ has degree at least 1. ∎
In this section, a matching will refer to a matching between agents and bundles (and not between agents and items as in previous sections). Thus, a matching is a mapping $\rho:A\to\mathcal{T}\cup\{\bot\}$ such that $\rho(i)=\rho(k)$ implies $\rho(i)=\rho(k)=\bot$. A perfect matching has $\rho(i)\neq\bot$ for every $i\in A$. Matchings may use pairs $(i,T_{k})$ that are not in $E$; we say that $\rho$ is a matching in the bipartite graph $\mathcal{K}=(A\cup\mathcal{T},E)$ if $(i,\rho(i))\in E$ whenever $\rho(i)\neq\{\bot\}$. For two matchings $\rho$ and $\tau$, an alternating path between $\rho$ and $\tau$ is a path $P=(i_{1},S_{i_{1}},i_{2},\ldots,S_{i_{k-1}},i_{\ell},S_{i_{\ell}})$ such that $\rho(i_{t})=S_{i_{t}}$, $t=1,\ldots,\ell$, $\tau(i_{t+1})=S_{i_{t}}$, $t=1,\ldots,\ell-1$. The following lemma is immediate from the definition of the $\nicefrac{{1}}{{2}}$-EFX feasibility graph.
Lemma 3.4.
If the $\nicefrac{{1}}{{2}}$-EFX feasibility graph $\mathcal{K}=(A\cup\mathcal{T},E)$ of an allocation $\mathcal{T}$ contains a perfect matching $\rho$, then $(i,\rho(i))_{i\in A}$ is a $\nicefrac{{1}}{{2}}$-EFX allocation.
Input: Partial allocation $\mathcal{T}$.
Output: Partial allocation $\mathcal{R}$ such that either $\operatorname*{NSW}(\mathcal{R})\geq\operatorname*{NSW}(\mathcal{T})$ and $\cup_{i}R_{i}\subsetneq\cup_{i}T_{i}$, or $\operatorname*{NSW}(\mathcal{R})\geq\tfrac{1}{2}\operatorname*{NSW}(\mathcal{T})$ and $\mathcal{R}$ is $\nicefrac{{1}}{{2}}$-EFX.
1 $\mathcal{S}\leftarrow\mathcal{T}$
2 $\rho(i)\leftarrow\bot$ for all $i\in A$
3 while $\rho$ is not a perfect matching in $\mathcal{K}$ do
 $\mathcal{K}=(A\cup\mathcal{S},E)\leftarrow\nicefrac{{1}}{{2}}$-EFX feasibility graph of $\mathcal{S}$
    // Definition 3.2
 $\mathcal{L}\leftarrow\{S_{i},i\in A\ |\ S_{i}\subsetneq T_{i}\}$
    // set of trimmed down bundles
   Define matching $\tau$ with $\tau(i)=S_{i}$ for all $i\in A$
    // candidate matching
 $\rho\leftarrow$ matching in $\mathcal{K}$ where
    // Lemma 3.5
4
  • $(a)$
    all bundles in $\mathcal{L}$ are matched,
  • $(b)$
    $|\{i:\rho(i)=S_{i}\}|$ is maximized subject to $(a)$, and
  • $(c)$
    $\rho$ is maximum subject to $(a)$ and $(b)$
5 if $\exists i_{1}\in A$ not matched in $\rho$ then
6    $(h,g_{h})\leftarrow\arg\max_{k\in A,g_{k}\in S_{k}}v_{i_{1}}(S_{k}-g_{k})$
7    if $v_{h}(S_{h}-g_{h})\geq\frac{1}{2}\cdot v_{h}(T_{h})$ then
8       $S_{h}\leftarrow S_{h}-g_{h}$
9    else
       $P=(i_{1},S_{i_{1}},i_{2},\dots,S_{i_{\ell-1}},i_{\ell},S_{i_{\ell}})$ $\leftarrow$ alternating path between $\tau$ and $\rho$ starting at $i_{1}$ and ending at either $S_{i_{\ell}}=S_{h}$ or an unmatched $S_{i_{\ell}}\neq S_{h}$
          // Lemma 3.6
10         Construct $\mathcal{R}$:
11            $R_{i_{1}}\leftarrow S_{h}-g_{h}$
12            for $f\leftarrow 2$ to $\ell$ do $R_{i_{f}}\leftarrow S_{i_{f-1}}$
13            for $i\in A\setminus(\{i_{1},\dots,i_{\ell}\}\cup\{h\})$ do $R_{i}\leftarrow T_{i}$
14       if $P$ ends at an unmatched bundle $S_{i_{\ell}}\neq S_{h}$ then
15               $R_{h}\leftarrow T_{h}\setminus(S_{h}-g_{h})$
16
17       return $\mathcal{R}$
18
19
20
21return $\mathcal{R}=(\rho(i))_{i\in A}$
Algorithm 3 MakeFairOrEfficient($\mathcal{T}$)
We now give an overview of Algorithm 3. For an input partial allocation $\mathcal{T}=(T_{i})_{i\in A}$, it returns a partial allocation $\mathcal{R}$ that satisfies one of the alternatives in Lemma 3.1: either (i) $\operatorname*{NSW}(\mathcal{R})\geq\operatorname*{NSW}(\mathcal{T})$ and $\cup_{i}R_{i}\subsetneq\cup_{i}T_{i}$, or (ii) $\operatorname*{NSW}(\mathcal{R})\geq\tfrac{1}{2}\operatorname*{NSW}(\mathcal{T})$ and $\mathcal{R}$ is $\nicefrac{{1}}{{2}}$-EFX.
The algorithm gradually ‘trims down’ the bundles $\mathcal{T}$. That is, we maintain a partial allocation $\mathcal{S}=(S_{i})_{i\in A}$ with $S_{i}\subseteq T_{i}$ throughout; these are initialized as $S_{i}=T_{i}$. Every main loop of the algorithm either terminates by constructing an allocation $\mathcal{R}$ satisfying (ii), or removes an item from one of the $S_{h}$ sets. The other possible termination option is when the $\nicefrac{{1}}{{2}}$-EFX feasibility graph of $\mathcal{S}$ contains a perfect matching $\rho$. In this case, we return $\mathcal{R}=(\rho(i))_{i\in A}$. This is a $\nicefrac{{1}}{{2}}$-EFX allocation by Lemma 3.4; Lemma 3.7 shows it also satisfies $\operatorname*{NSW}(\mathcal{R})\geq\tfrac{1}{2}\operatorname*{NSW}(\mathcal{T})$ and is thus a suitable output of type (ii).
At the beginning of each main loop, we define two matchings. The first is the perfect matching $\tau$ that simply defines $\tau(i)=S_{i}$ for all $i\in A$. The second is a matching $\rho$ in $\mathcal{K}$. This is required to satisfy three properties: First, it matches all trimmed down bundles, i.e., all bundles $S_{i}$ with $S_{i}\subsetneq T_{i}$. Second, $|\{i:\rho(i)=S_{i}\}|$ is maximized subject to the first requirement. Third, subject to these requirements, $\rho$ is chosen as a maximal matching. (The existence of such a matching is guaranteed by Lemma 3.5 below).
If $\rho$ is not perfect, then we consider an unmatched agent $i_{1}$, and find the bundle that maximizes $i_{1}$’s utility after removal of one item. Let $(S_{h},g_{h})\in\arg\max_{k\in A,g\in S_{k}}v_{i_{1}}(S_{k}-g)$. If agent $h$’s value of $S_{h}-g_{h}$ is at least $\frac{1}{2}$ times their value for the original bundle $T_{h}$, then we remove $g_{h}$ from $S_{h}$ and the main loop finishes. Otherwise, we construct an alternating path between $\rho$ and $\tau$, denoted as $P=(i_{1},S_{i_{1}},i_{2},S_{i_{2}},\dots,S_{i_{\ell-1}},i_{\ell},S_{i_{\ell}})$, starting with $i_{1}$ and ending with either $S_{i_{\ell}}=S_{h}$ or an unmatched bundle $S_{i_{\ell}}\neq S_{h}$. Lemma 3.6 shows that such a $P$ exists. Using $P$, we construct an allocation $\mathcal{R}$ in line 3. Lemma 3.7 shows that this is a suitable output of type (i).
3.1.1 Analysis
The number of iterations of the repeat loop is at most $m$, because in each iteration except the final one, an item is removed from one of the bundles. Noting a maximum matching in line 3 and alternating path in line 3 can be found in strongly polynomial-time, Algorithm 3 runs in strongly polynomial-time.
The next lemma guarantees that the matching $\rho$ is well-defined. The proof follows similarly as in [18].
Lemma 3.5.
In each iteration of the repeat loop in Algorithm 3, a matching exists in $\mathcal{K}$ where all bundles of $\mathcal{L}$ are matched.
Proof.
Let $E^{(t)}$ denote the edge set of the $\nicefrac{{1}}{{2}}$-EFX feasibility graph and $\mathcal{L}^{(t)}$ the set of trimmed down bundles, and $\rho^{(t)}$ the maximum matching in the $t$-th iteration.
We show by induction that there exists a matching $\rho^{(t)}$ such that all bundles in $\mathcal{L}^{(t)}$ are matched. At the beginning of the first iteration, $\mathcal{L}^{(1)}$ is empty, so the claim is clearly true. Suppose the claim is true until the beginning of $(t+1)$-st iteration. Let $\mathcal{S}$ denote the trimmed down bundles in the $t$-th iteration, and let $i_{1}$ be the unmatched agent, and $(S_{h},g_{h})$ the bundle and item selected in line 3.
By the requirement that $|\{i:\rho(i)=S_{i}\}|$ is maximized subject to all trimmed down bundles being matched, we have $(i_{1},S_{i_{1}})\not\in E^{(t)}$. This means that $2v_{i_{1}}(S_{i_{1}})<\max_{\ell\in A,j\in S_{\ell}}v_{i_{1}}(S_{\ell}-j)=v_{i_{1}}(S_{h}-g_{h})$ by the choice of $h$. Thus, $(i_{1},S_{h})\in E^{(t)}$ follows.
Note that $\mathcal{L}^{(t+1)}=\mathcal{L}^{(t)}\cup\{h\}$. Consider the $\nicefrac{{1}}{{2}}$-EFX feasibility graph in the $(t+1)$-st iteration. Since all bundles different from $S^{\prime}_{h}:=S_{h}-g_{h}$ remained unchanged, for every edge $(i,S_{k})\in E^{(t)}$ with $k\neq h$ it follows that $(i,S_{k})\in E^{(t+1)}$. According to Definition 3.2, $(i_{1},S_{h}^{\prime})\in E^{(t+1)}$. Let us define $\rho^{\prime}$ as
$\rho^{\prime}(i):=\begin{cases}S^{\prime}_{h}&\mbox{if }i=i_{1}\,,\\ \rho^{(t)}(i)&\mbox{if }i\neq i_{1}\mbox{ and }\rho^{(t)}(i)\neq S_{h}\,,\\ \bot&\mbox{otherwise}\,.\end{cases}$
By the above, this gives a matching in $E^{(t+1)}$, and it matches all bundles in $\mathcal{L}^{(t+1)}=\mathcal{L}^{(t)}\cup\{h\}$. ∎
Lemma 3.6.
The alternating path $P$, as described in line 3 of Algorithm 3, exists.
Proof.
Since $i_{1}$ is an unmatched agent and the requirement that $|\{i:\rho(i)=S_{i}\}|$ is maximized subject to all trimmed down bundles being matched in the maximum matching $\rho$ in line 3, we must have $(i_{1},S_{i_{1}})\not\in E$. If $\rho(i_{2})=S_{i_{1}}$ for an agent $i_{2}\in A$, then we continue with $S_{i_{2}}$, otherwise we stop. Continuing this way, we eventually reach either $S_{i_{\ell}}=S_{h}$ or an unmatched bundle $S_{i_{\ell}}\neq S_{h}$. ∎
Lemma 3.7.
If Algorithm 3 returns an allocation $\mathcal{R}$ in line 3, then $\operatorname*{NSW}(\mathcal{R})\geq\tfrac{1}{2}\operatorname*{NSW}(\mathcal{T})$ and $\mathcal{R}$ is $\nicefrac{{1}}{{2}}$-EFX. If it returns $\mathcal{R}$ in line 3, then $\operatorname*{NSW}(\mathcal{R})\geq\operatorname*{NSW}(\mathcal{T})$ and $\cup_{i}R_{i}\subsetneq\cup_{i}T_{i}$.
Proof.
Let us start with the case when a perfect matching $\mathcal{R}=(\rho(i))_{i\in A}$ is returned in line 3. The $\nicefrac{{1}}{{2}}$-EFX property follows by Lemma 3.4. Let us show $\operatorname*{NSW}(\mathcal{R})\geq\tfrac{1}{2}\operatorname*{NSW}(\mathcal{T})$.
Throughout the algorithm, $v_{i}(S_{i})\geq\frac{1}{2}v_{i}(T_{i})$ is maintained according to the condition on bundle trimming. By the definition of the $\nicefrac{{1}}{{2}}$-EFX feasibility graph, either $R_{i}=S_{i}$, or $v_{i}(R_{i})>2v_{i}(S_{i})$. Therefore, we have
$\forall i$ $:v_{i}(R_{i})\geq v_{i}(S_{i})\geq\frac{1}{2}v_{i}(T_{i})\,.$ (7)
Consequently,
$\operatorname*{NSW}(\mathcal{R})=\left(\prod_{i\in A}v_{i}(R_{i})\right)^{1/n}\geq\frac{1}{2}\left(\prod_{i\in A}v_{i}(T_{i})\right)^{1/n}\geq\tfrac{1}{2}\operatorname*{NSW}(\mathcal{T})\,.$
Consider now the case when the algorithm terminated with $\mathcal{R}$ in line 3. We need to show $\operatorname*{NSW}(\mathcal{R})\geq\operatorname*{NSW}(\mathcal{T})$ and $\cup_{i}R_{i}\subsetneq\cup_{i}T_{i}$. We distinguish two cases based on whether $S_{i_{\ell}}=S_{h}$, or whether $S_{i_{\ell}}(\neq S_{h})$ is an unmatched bundle. For the first case, we have
$v_{i_{f}}(R_{i_{f}})$ $=v_{i_{f}}(S_{i_{f-1}})>2v_{i_{f}}(S_{i_{f}})\geq v_{i_{f}}(T_{i_{f}}),\ \forall f\in\{2,\dots,\ell\}\,.$
The first inequality follows from the definition of the $\nicefrac{{1}}{{2}}$-EFX feasibility graph, and the second inequality follows from the bundle trimming condition maintained throughout the algorithm. Similarly, we have
$v_{i_{1}}(R_{i_{1}})$ $=v_{i_{1}}(S_{h}-g_{h})>2v_{i_{1}}(S_{i_{1}})\geq v_{i_{1}}(T_{i_{1}})\,.$
This implies
$\operatorname*{NSW}(\mathcal{R})=\left(\prod_{i\in A}v_{i}(R_{i})\right)^{1/n}>\left(\prod_{i\in A}v_{i}(T_{i})\right)^{1/n}=\operatorname*{NSW}(\mathcal{T})\,.$
Since we do not assign $g_{h}$ to any agent in $\mathcal{R}$, we must have $\cup_{i}R_{i}\subsetneq\cup_{i}T_{i}$.
In the second case, since $S_{i_{\ell}}$ is an unmatched bundle in $\rho$ by the choice of the path $P$, we have $S_{i_{\ell}}\notin\mathcal{L}$ by the requirements on $\rho$. That is, $S_{i_{\ell}}=T_{i_{\ell}}$. By the definition of the $\nicefrac{{1}}{{2}}$-EFX feasibility graph, we have
$v_{i_{\ell}}(R_{i_{\ell}})$ $=v_{i_{\ell}}(S_{i_{\ell-1}})>2v_{i_{\ell}}(S_{i_{\ell}})=2v_{i_{\ell}}(T_{i_{\ell}})\,$, (8)
$v_{i_{f}}(R_{i_{f}})$ $=v_{i_{f}}(S_{i_{f-1}})>2v_{i_{f}}(S_{i_{f}})\geq v_{i_{f}}(T_{i_{f}}),\ \forall f\in\{2,\dots,\ell-1\}\,$,
$v_{i_{1}}(R_{i_{1}})$ $=v_{i_{1}}(S_{h}-g_{h})>2v_{i_{1}}(S_{i_{1}})\geq v_{i_{1}}(T_{i_{1}})$
$v_{h}(R_{h})$ $=v_{h}(T_{h}\setminus(S_{h}-g_{h}))>\tfrac{1}{2}v_{h}(T_{h})\,.$
The last inequality follows from subadditivity using $v_{h}(T_{h})\leq v_{h}(S_{h}-g_{h})+v_{h}(T_{h}\setminus(S_{h}-g_{h}))\leq\frac{1}{2}v_{h}(T_{h})+v_{h}(T_{h}\setminus(S_{h}-g_{h}))$. Using (8), we get
$\operatorname*{NSW}(\mathcal{R})=\left(\prod_{i\in A}v_{i}(R_{i})\right)^{1/n}>\left(\prod_{i\in A}v_{i}(T_{i})\right)^{1/n}=\operatorname*{NSW}(\mathcal{T})\,.$
Finally, since we do not assign items in $T_{i_{\ell}}$ to any agent in $\mathcal{R}$, we must have $\cup_{i}R_{i}\subsetneq\cup_{i}T_{i}$. Note that if $T_{i_{\ell}}=\emptyset$, then $\operatorname*{NSW}(\mathcal{T})=0$ and $R_{i}=\emptyset,\forall i$ is a suitable output of type (ii). ∎
3.2 Completing the partial allocation
In this Section, we derive Theorem 1.2 from Lemma 3.1. The corresponding algorithmic proof is given in Algorithm 4, which relies on two subroutines MakeFairOrEfficient (Algorithm 3) and EnvyFreeCycle, the latter being an envy cycle procedure adapted from [48] and described below.
The input of Algorithm 4 is an allocation $\mathcal{S}$ that is $\alpha$-approximation to the symmetric NSW problem. It then repeatedly calls Algorithm 3 until the final allocation is $\nicefrac{{1}}{{2}}$-EFX and $2\alpha$-approximation to the symmetric NSW problem. Recall that the output of this subroutine is either a partial allocation $\mathcal{T}^{\prime}$ that satisfies either $\operatorname*{NSW}(\mathcal{T}^{\prime})\geq\operatorname*{NSW}(\mathcal{T})$ and $\cup_{i}T_{i}^{\prime}\subsetneq\cup_{i}T_{i}$, or $\operatorname*{NSW}(\mathcal{T}^{\prime})\geq\nicefrac{{1}}{{2}}\operatorname*{NSW}(\mathcal{T})$ and $\mathcal{T}^{\prime}$ is $\nicefrac{{1}}{{2}}$-EFX. Since $\cup_{i}T_{i}^{\prime}\subsetneq\cup_{i}T_{i}$ in each call in the first case, the number of calls to Algorithm 3 is at most $m$.
Input: Allocation $\mathcal{S}$ that is $\alpha$-approximation to the NSW problem $(A,G,(v_{i})_{i\in A})$.
Output: Allocation $\mathcal{T}$ that is $\nicefrac{{1}}{{2}}$-EFX and $2\alpha$-approximation to the symmetric NSW problem.
1 $\mathcal{T}\leftarrow\mathcal{S}$
2 while $\mathcal{T}$ is not $\nicefrac{{1}}{{2}}$-EFX do
 $\mathcal{T}\leftarrow\textnormal{{MakeFairOrEfficient}}(\mathcal{T})$
    // Algorithm 3
3
$U\leftarrow G\setminus\cup_{i}T_{i}$
  // set of unallocated items
4 while $\exists i\in A\,,\exists j\in U\,:\,v_{i}(T_{i})<v_{i}(j)$ do
5   Let $j\in U$ be such that $v_{i}(T_{i})<v_{i}(j)$ for some agent $i$
6 $U\leftarrow(U\cup T_{i})-j$
7 $T_{i}\leftarrow\{j\}$
8
[…]
10 return $\mathcal{T}$
Algorithm 4 Guaranteeing $\nicefrac{{1}}{{2}}$-EFX for the symmetric NSW problem
At this point, we have a $\nicefrac{{1}}{{2}}$-EFX partial allocation $\mathcal{T}$ with $\operatorname*{NSW}(\mathcal{T})\geq\tfrac{1}{2}\operatorname*{NSW}(\mathcal{S})$. The rest of Algorithm 4 allocates the remaining items $U=G\setminus(\cup_{i\in A}T_{i})$ so that $\operatorname*{NSW}(\mathcal{T})$ does not decrease, and the $\nicefrac{{1}}{{2}}$-EFX property is maintained.
First, we modify the allocation in the second repeat loop to ensure that each agent’s value for their bundle is at least their value for each remaining item in $U$. This is done by swapping an agent’s bundle $T_{i}$ with a singleton item $j\in U$ whenever $i$ values $j$ more than the entire bundle $T_{i}$.
Finally, we run the envy-cycle procedure EnvyFreeCycle($\mathcal{T},U$), introduced in [48], to allocate the remaining unassigned items in $U$, starting from the partial allocation $\mathcal{T}=(T_{i})_{i\in A}$. The procedure maintains and repeatedly updates a directed (envy) graph $D=(A,E)$, where each node corresponds to an agent, and there is a directed edge $(i,j)\in E$ if agent $i$ envies agent $j$, i.e., $v_{i}(T_{i})<v_{i}(T_{j})$.
The procedure proceeds as follows. If the graph $D$ contains a directed cycle, then the bundles of the agents along the cycle are rotated: each agent receives the bundle of the next agent in the cycle. This rotation strictly improves the utility of each agent involved in the cycle, while leaving the utilities of all other agents unchanged. If the graph contains no cycles, then it must have at least one source agent—an agent who is not envied by any other agent. In this case, the procedure selects an arbitrary item from $U$ and assigns it to a source agent. The envy graph is then updated to reflect the new allocation. The procedure continues until all items in $U$ have been assigned. We refer the interested reader to [48] for more details.
We now verify the correctness and efficiency of this algorithm.
Lemma 3.8.
The second repeat loop of Algorithm 4 is repeated at most $nm$ times. It maintains the $\nicefrac{{1}}{{2}}$-EFX property and $\operatorname*{NSW}(\mathcal{T})$ is non-decreasing.
Proof.
The bound on the number of swaps follows because every agent $i\in A$ may swap their bundle at most $m$ times. After the first swap, they maintain a singleton bundle, and they can swap their bundle for the same item $j$ only once, since their valuation $v_{i}(T_{i})$ strictly increases in each swap.
It is immediate that $\operatorname*{NSW}(\mathcal{T})$ is non-decreasing. It is left to show that the $\nicefrac{{1}}{{2}}$-EFX property is maintained. Let $i\in A$ be the agent who swapped their bundle $T_{i}$ for $T^{\prime}_{i}=\{j\}$ in the current iteration. Then, the value of $i$’s own bundle increased while the allocation of everyone else remained the same. Hence, agent $i$ cannot violate the $\nicefrac{{1}}{{2}}$-EFX property. For the other agents $k\neq i$, $v_{k}(T_{k})\geq\nicefrac{{1}}{{2}}\cdot v_{k}(T_{i}^{\prime}-g)$ for all $g\in T_{i}^{\prime}$ trivially holds, since $T_{i}^{\prime}$ is a singleton. ∎
Property (9) below is satisfied after the second repeat loop. Hence, the next lemma completes the analysis of Algorithm 4.
Lemma 3.9.
The subroutine EnvyFreeCycle($\mathcal{T},U$) terminates in $O(n^{3}m)$ time, and $\operatorname*{NSW}(\mathcal{T})$ is non-decreasing. Assume that $\mathcal{T}=(T_{i})_{i\in A}$ is $\nicefrac{{1}}{{2}}$-EFX, and
$v_{i}(T_{i})\geq v_{i}(j)\quad\forall i\in A,\forall j\in U\,.$ (9)
Then, EnvyFreeCycle($\mathcal{T},U$) also maintains the $\nicefrac{{1}}{{2}}$-EFX property.
Proof.
The running time analysis is the same as in [48]. Finding and removing a cycle in the envy-graph can be done in $O(n^{2})$ time. Further, whenever swapping around a cycle, at least one edge is removed from the envy graph. New edges can only be added when we allocate new items from $U$, with at most $n$ edges every time. Since $|U|\leq m$, the total number of new edges added throughout is $nm$. This yields the overall $O(n^{3}m)$ bound.
Again, it is immediate that $\operatorname*{NSW}(\mathcal{T})$ is non-decreasing in every step. We need to show that the $\nicefrac{{1}}{{2}}$-EFX property is maintained both when swapping around cycles and when adding new items from $U$. When swapping around a cycle, this follows since the set of bundles remains the same, and no agent’s value decreases.
Consider the case when a source agent say $i$, gets a new item $j$: their new bundle becomes $T^{\prime}_{i}=T_{i}+j$. Note that $i$ is the only agent whose bundle grows; all other bundles remain the same. For any $k\neq i$ and any $g\in T^{\prime}_{i}$, we can bound
$v_{k}(T^{\prime}_{i}-g)=v_{k}(T_{i}+j-g)\leq v_{k}(T_{i})+v_{k}(j)\leq 2v_{k}(T_{k})\,$,
showing that the $\nicefrac{{1}}{{2}}$-EFX property is preserved. Here, the first inequality follows by subadditivity and monotonicity. The second inequality uses (9), and that $v_{k}(T_{k})\geq v_{k}(T_{i})$, since $i$ was a source node in the envy graph. ∎
4 Conclusion
We have shown a $(4+\varepsilon)$-approximation algorithm for the symmetric NSW problem with submodular valuations. Moreover, our algorithm yields an $\exp\big{(}\log(3n)+\sum_{i=1}^{n}w_{i}\log w_{i}\big{)}\cdot(\mathrm{e}+\varepsilon)$-approximation algorithm for the asymmetric NSW problem under submodular valuations (see Remark 2.15, [52]). Subsequent work [10, 31, 32] has significantly extended these results, leading to a small constant factor approximation independent of the weights. However, there are still several directions and open problems to explore. An obvious one is to improve the approximation ratio for the symmetric case. The current hardness of approximation stands at $\frac{\mathrm{e}}{\mathrm{e}-1}\simeq 1.58$ for submodular valuations, which is the same as the optimal factor for maximizing utilitarian social welfare. It would be interesting to prove a separation between the two optimization objectives for submodular valuations.
There are several open questions on the existence of EFX and its relaxations for submodular valuations. A significant one is: Does a (complete) $\alpha$-EFX allocation exist for $\alpha>1/2$? Here, even without imposing any efficiency requirements.
Acknowledgements
We are grateful to Tomasz Ponitka for providing an improved version of Lemma 2.9 (previous version had factor $6$ instead of $4$) and for pointing out a couple of minor issues in Section 3 of the previous version that have been corrected in this paper. We also thank Bento Natura for the improved bound included in Remark 2.15.
Appendix A A General Rematching Lemma
We now discuss the general statement on matchings underlying our Rematching Lemma (Lemma 2.1). We first show a simple and natural statement on bipartite graphs. We were unable to find a previous explicit occurence in the literature. For 0 and -$\infty$ edge weights, this was shown by Brualdi and Scrigmer [16], and by Brualdi and Mason [15]. This special case implies that the set system representing a transversal matroid of rank $d$ can be reduced to a representation of size $d$. Lemma A.1 implies the analogous statement for valuated transversal matroids and has been used implicitly in the literature.
Lemma A.1.
Let $\mathcal{G}=(A,B;c)$ be a complete bipartite graph with edge weights $c\in({\mathbb{R}}\cup\{-\infty\})^{A\times B}$. Let $\tau\,:\,A\to B$ be a maximum-weight $A$-perfect matching, and let $H\coloneqq\tau(A)\subseteq B$ denote the set of matched nodes in $B$. Then, for any $X\subseteq A$, there is a maximum-weight $X$-perfect matching $\rho\,:\,X\to B$ such that $\rho(X)\subseteq H$.
Proof.
Take any maximum-weight $X$-perfect matching $\rho\,:\,X\to B$ with minimum size of $\rho(X)\setminus H$. For the sake of contradiction, assume $\rho(X)\setminus H\neq\emptyset$ and let $j\in\rho(X)\setminus H$. Let us consider the symmetric difference $\tau\Delta\rho$; this consists of cycles and paths. Since $\not\in H, jj$ is an endpoint of an alternating path $P$ between $\rho$ and $\tau$ that starts in $j$ with a $\rho$-edge and ends in $H$ with a $\tau$-edge (since $\tau$ is an $A$-perfect matching). Let $Y\subseteq A$ be the set of nodes incident to $P$. Clearly, $Y\subseteq X$. We get another $A$-perfect matching $\tau^{\prime}\,:\,A\to B$ by flipping $\tau$ along $P$, i.e., setting $\tau^{\prime}(i)=\rho(i)$ for $i\in Y$ and $\tau^{\prime}(i)=\tau(i)$ for $i\in A\setminus Y$. Similarly, we obtain another $X$-perfect matching $\rho^{\prime}\,:\,X\to B$ by flipping $\rho$ along $P$, i.e., setting $\rho^{\prime}(i)=\tau(i)$ for $i\in Y$ and $\rho^{\prime}(i)=\rho(i)$ for $i\in X\setminus Y$.
By the optimality of $\tau$, the weight of $\tau$-edges in $P$ is at least the weight of $\rho$-edges in $P$. This implies that the weight of $\rho^{\prime}$ is at least the weight of $\rho$. This contradicts the choice of $\rho$ since $\rho^{\prime}(X)\setminus H=(\rho(X)\setminus H)\setminus\{j\}$. ∎
Lemma 2.1 requires a slight extension of the above statement, namely when we may create multiple copies of the nodes in $B\setminus H$. The proof is similar as above.
Lemma A.2.
Let $\mathcal{G}=(A,B;c)$ be a complete bipartite graph with edge weights $c\in({\mathbb{R}}\cup\{-\infty\})^{A\times B}$. Let $\tau\,:\,A\to B$ be a maximum-weight $A$-perfect matching, and let $H\coloneqq\tau(A)\subseteq B$ denote the set of matched nodes in $B$. For some integer $k\geq 1$, let us obtain from $\mathcal{G}$ the bipartite graph $\mathcal{G}^{\prime}=(A,B^{\prime};c^{\prime})$ by creating $k$ copies of every node in $B\setminus H$; thus, $|B^{\prime}|=k|B|-(k-1)|A|$. Then, for any $X\subseteq A$, there is a maximum-weight $X$-perfect matching $\rho\,:\,X\to B^{\prime}$ such that $\rho(X)\subseteq H$.
Proof.
The case $k=1$ is identical to Lemma A.1. The proof of $k>1$ follows similarly: we consider the symmetric difference of $\tau$ (naturally embedded into the larger graph $\mathcal{G}^{\prime}$) and $\rho$. We take an alternating path $P$ starting from a node $j^{\prime}\in\rho(X)\setminus H$. Assume $j^{\prime}$ is one of the $k$ copies of a node $j\in B\setminus H$. The $A$-perfect matching $\tau^{\prime}\,:\,A\to B^{\prime}$ obtained from $\tau$ by flipping along $P$ will assign $\tau^{\prime}(i^{\prime})=j^{\prime}$ for some $i^{\prime}\in A$. This naturally maps back to an $A$-perfect matching $\bar{\tau}\,:\,A\to B$ of the same weight in the original graph, by setting $\bar{\tau}(i^{\prime})=j$, and $\bar{\tau}(k)=\tau^{\prime}(k)$ for $k\neq i^{\prime}$. This again shows that the weight of $\tau$-edges in $P$ is at least the weight of $\rho$-edges in $P$, and we obtain a contradiction by constructing a better $X$-perfect matching $\rho^{\prime}\,:\,A\to B^{\prime}$ as in the $k=1$ case. ∎